40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 14 09:28:46 2021, Yehonathan, Update, PSL, PMC unlocked 
    Reply  Thu Oct 14 11:25:49 2021, Yehonathan, Update, PSL, PMC unlocked BS_OSEM_Sensor_PSD.pdfBS_OSEM_Sensor_PSD_AfterReconnectingCables.pdf
Message ID: 16401     Entry time: Thu Oct 14 11:25:49 2021     In reply to: 16400
Author: Yehonathan 
Type: Update 
Category: PSL 
Subject: PMC unlocked 

{Yehonathan, Anchal}

I went to get a sandwich around 10:20 AM and when I came back BS was moving like crazy. We shutdown the watchdog.

We look at the spectra of the OSEMs (attachment 1). Clearly, the UR sensing is bad.

We took the BS sattelite box out. Anchal opened the box and nothing seemed wrong visually. We returned the box and connected it to the fake OSEM box. The sensor spectra seemed normal.

We connected the box to the vacuum chamber and the spectra is still normal (attachment 2).

We turn on the coils and the motion got damped very quickly (RMS <0.5mV).

Either the problem was solved by disconnecting and connecting the cables or it will come back to haunt us.

 

 

 

Attachment 1: BS_OSEM_Sensor_PSD.pdf  19 kB  Uploaded Thu Oct 14 12:29:54 2021  | Hide | Hide all
BS_OSEM_Sensor_PSD.pdf
Attachment 2: BS_OSEM_Sensor_PSD_AfterReconnectingCables.pdf  35 kB  Uploaded Thu Oct 14 13:10:46 2021  | Hide | Hide all
BS_OSEM_Sensor_PSD_AfterReconnectingCables.pdf
ELOG V3.1.3-